Activity Predecessor Optimistic Time (a) Most Probable Time (m) Pessimistic Time (b)
A None 2 6 10
B None 2 4 6
C A 2 4 6
D A 4 6 8
E B 6 8 22
F B 6 8 10
G C,D 2 8 12
H E,F 2 7 12
I G,H 4 8 16

Question for Blank 1: What is the Critical Path?

Question for Blank 2: What is the Expected Duration of the Project? round to nearest whole number

Question for Blank 3: If a Target of 35 days is given, what is the probability of the project being completed on or before the Target date?
Answer #1

EXPECTED TIME

FORMULA:

(a + (4m) + b) / 6

A = (2 + (4 * 6) + 10) / 6 = 6

B = (2 + (4 * 4) + 6) / 6 = 4

C = (2 + (4 * 4) + 6) / 6 = 4

D = (4 + (4 * 6) + 8) / 6 = 6

E = (6 + (4 * 8) + 22) / 6 = 10

F = (6 + (4 * 8) + 10) / 6 = 8

G = (2 + (4 * 8) + 12) / 6 = 7.67

H = (2 + (4 * 7) + 12) / 6 = 7

I = (4 + (4 * 8) + 16) / 6 = 8.67


VARIANCE

FORMULA:

((b - a) / 6) ^ 2

A = (10 - 2) / 6) ^ 2 = 1.78

B = (6 - 2) / 6) ^ 2 = 0.44

C = (6 - 2) / 6) ^ 2 = 0.44

D = (8 - 4) / 6) ^ 2 = 0.44

E = (22 - 6) / 6) ^ 2 = 7.11

F = (10 - 6) / 6) ^ 2 = 0.44

G = (12 - 2) / 6) ^ 2 = 2.78

H = (12 - 2) / 6) ^ 2 = 2.78

I = (16 - 4) / 6) ^ 2 = 4


CPM ANALYSIS


ACTIVITY

DURATION

ES

EF

LS

LF

SLACK

A

6

0

6

1.33

7.33

1.33

B

4

0

4

0

4

0

C

4

6

10

9.33

13.33

3.33

D

6

6

12

7.33

13.33

1.33

E

10

4

14

4

14

0

F

8

4

12

6

14

2

G

7.67

12

19.67

13.33

21

1.33

H

7

14

21

14

21

0

I

8.67

21

29.67

21

29.67

0



FORWARD PASS = ES = MAX EF OF THE PREDECESSOR

EF = ES + DURATION

BACKWARD PASS = LF = MIN LS OF ALL PREDECESSORS

LS = LF - DURATION

CRITICAL PATH = PATH WITH LONGEST DURATION AND 0 NET SLACK

CRITICAL PATH & DURATION

B + E + H + I = 29.67


WE CAN TAKE THE TIME OF COMPLETION TO BE EITHER 29.67 OR 30, TO ROUND IT UP TO THE NEAREST INTEGER

VARIANCE OF CRITICAL PATH = 0.44 + 7.11 + 2.78 + 4 = 14.33

Due time = 35

Expected time = 30

Stdev = sqrt(variance) = sqrt(14.33) = 3.7855(ROUNDED)

Z = DUE - EXPECTED / STDEV

Z = 35 - 30 / 3.7855 = 1.3208

Z value of 1.3208 gives a distribution of 0.9067 OR 90.67 % probability of completion of the project by the due time of 35

**INTERMEDATECALCULATIONS FOR PROBABILISTIC ESTIMATES ROUNDED TO 4 DECIMAL PLACES.